# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v1

@@ Abstract
TIPO (Text to Image with text pre-sampling for Prompt Optimization) is an innovative framework designed to enhance text-to-image (T2I) generation by language model (LM) for automatic prompt engineering. By refining and extending user-provided prompts, TIPO bridges the gap between simple inputs and the detailed prompts required for high-quality image generation.
Unlike previous approaches that rely on Large Language Models (LLMs) or reinforcement learning (RL), TIPO adjusts user input prompts with the distribution of a trained prompt dataset, eliminating the need for complex runtime cost via lightweight model. This pre-sampling approach enables efficient and scalable prompt optimization. Experimental results demonstrate TIPO’s effectiveness in improving aesthetic scores, reducing image corruption, and better aligning generated images with dataset distributions. These findings highlight the critical role of prompt engineering in T2I systems and open avenues for broader applications of automatic prompt refinement.
@@ 1 Introduction
Recent advancements in Text-to-Image (T2I) generative models have revolutionized creative applications [36, 6, 19, 41, 38, 39, 43, 40, 37, 42, 12, 11, 28, 20, 7]. These models tend to perform better when provided with longer, more detailed prompts, which describes specific elements like style, composition, and context. However, this reliance on highly descriptive inputs is a significant bottleneck for generating high-quality images, especially for users who prefer or require simpler inputs. The need for intricate, nuanced prompts often results in a higher barrier to entry, limiting accessibility and ease of use for those unfamiliar with the complexities of prompt engineering. The previous works [31] tried to utilize LLMs to lenghthen prompt to solve tendency, however many ignore the root cause of why some prompts result in worse images and rely heavily on manually engineered prompts, which require substantial resources.
@@
Unlike previous approaches that focus on extending prompts with manually collected prompt sets [4, 15, 45, 21], directly use LLM to modify the prompts [26, 51] or utilize reinforcement learning like method which update model’s input or model itself from text-to-image feedback [22, 13, 31] without addressing underlying issues, TIPO introduces a method that optimizes prompts efficiently as shown Section 4.
@@
Figure 1: Comparison of prompt optimization methods. (a) uses instructions for prompt refinement but lacks understanding of T2I-specific prompts. (b) improves detail but limits variety. (c) optimizes for model feedback but are computationally expensive. (d) TIPO (ours) aligns prompts with the trained distribution of T2I models, ensuring prompt detail, consistency with training data, and controlled variety within model scope.
This paper presents TIPO’s theoretical foundations, implementation details, and experimental results demonstrating its effectiveness in improving aesthetic scores [17], reducing image corruption [32, 33], and better aligning generated images with dataset distributions [24, 35]. Our findings highlight the critical role of prompt optimization in advancing T2I technology and open avenues for broader applications of automatic prompt refinement.
@@ 2.3 Image Output-Based Tuning
Figure 2: Comparison of prompt optimization methods for a T2I model [7]. (a) yields a basic image with limited detail. (b) provides specific output but requires manual engineering. (c) introduce nonsensical elements (red boxes) in the teal region, where the model’s actual output distribution for a given prompt (green) and all the image distribution (blue) overlap outside the user’s expected output distribution (red). (d) captures user intent accurately, balancing detail and diversity.
@@ 3.1 Motivation and Challenges
Nowadays text-conditioned models [7, 37, 19] generate outputs based on user-provided text inputs. However, when users seek high-quality outputs or wish to explore diverse outputs while preserving key aspects of their input, they often need to manually expand their prompts. This manual prompt expansion can be time-consuming and labor-intensive. Moreover, simply creating a longer prompt does not guarantee improved results; it must be crafted in line with the model’s specific prompting conventions and interpretive nuances.
To address these challenges, previous works have aimed to automate the prompt optimization process. Some methods attempt to expand prompts using the model’s familiar vocabulary and stylistic tendencies [6], while others rely on direct feedback from the model to iteratively improve prompt alignment [22]. However, these approaches have limitations: if the structure or language does not fully align with the model’s prompting conventions, performance may degrade, and feedback-based approaches often involve significant computational costs.
Therefore, effective prompt optimization for text-conditioned models requires accurately capturing user intent and expanding the prompt in a way that is both precise and aligned with the model’s conventions for interpreting text inputs. This need for precision and alignment motivates a systematic approach to prompt formulation, which we define in the following section.
@@ 3.2 Formulation
The simplest approach to extend a simple prompt is by adding random words. However, this approach has limitations, as it often introduces internally conflicting information that results in out-of-distribution (OOD) prompt inputs, leading the model to generate nonsensical images. In such cases, although each added word may be meaningful in isolation, these random combinations can create prompts that are internally contradictory (and are not simply creative combinations). Thus, while random word extension typically yields unreliable results, it may occasionally produce a usable prompt by chance. To represent this random extension method, let $p_{r_{i}}$ denote the $i^{th}$ prompt created by adding random words to $p_{s}$. The resulting union of distributions fails to approximate the desired distribution.
@@
This scenario is illustrated in Figure 2, where “Simple Prompt + Random Words” shows a scattered distribution extending beyond the bounds of the desired distribution or even the “image” distribution itself.
4 TIPO Framework
4.1 Overview
TIPO introduces a dataset-driven approach to prompt optimization by training directly on the existing dataset. This avoids reliance on manually engineered prompts [4, 45, 15] or on computationally intensive RL techniques [22, 31]. Methods using LLMs rely on general language knowledge, which may misalign with specific T2I model needs [6, 51]. In contrast to previous approaches, TIPO captures the model’s training data distribution closely.
@@ 4.2 Distribution-Aligned Prompt Extension
where $p_{t_{i}}$ denotes the $i^{th}$ TIPO-generated prompt derived from $p_{s}$, and $\mathcal{I}_{p_{s}}$ represents the target distribution of image outputs associated with $p_{s}$. Through this systematic prompt refinement, Figure 2 “Simple Prompt + TIPO” illustrates the broader image space exploration enabled by TIPO while retaining relevance to the original prompt, effectively balancing diversity with coherence.
@@ 4.5 An Example Scenario and Workflow
Figure 3: (a) A generated image and prompts. (b) Prompt optimization of TIPO, starting from user input (short tag prompt $p_{s}$) to final output (detail prompt $p_{d}$). The shading from gray to light sky blue represents an increase in context richness in the prompt.
@@
$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\verb|<meta>|=\emptyset{}\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, ...}\right\}\\ \end{array}$
@@
    short_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed prompt.
@@
This example highlights how TIPO leverages an autoregressive LM to perform prompt extension in a stepwise, sequential processes, through short_to_tag, tag_to_long, and short_to_tag_to_long tasks within a single generation cycle. More detailed task definitions for TIPO can be found in Section 5.2. By iteratively refining prompts in this structured way, TIPO addresses the limitations of simplistic or randomly extended prompts, enabling it to generate high-quality, diverse images that closely align with user intent.
4.6 Versatile and Scalable Design of TIPO
The example in Section 4.5 demonstrates how TIPO’s structured approach enables it to create coherent, intent-aligned prompts. Building on this foundation, TIPO’s design extends smoothly to a wide range of models and datasets, showcasing its versatility and scalability.
A single TIPO model trained on dataset D can be applied to all text-to-image models trained on the same dataset, enabling broad applicability across different architectures and versions that share the same training data. Similarly, a TIPO model trained using caption method M (e.g. tag, NL captions) can be used with any T2I model trained with captions from method M. Experimental results in Section 6 further illustrate this adaptability, as TIPO successfully applies across multiple model architectures and training configurations, confirming its broad applicability and robustness.
5 Training TIPO model
In this section, we detail how to build the TIPO model based on the TIPO framework. For various T2I models [36, 6, 19, 41, 38, 39, 43, 40, 37, 42, 12, 11, 28, 20, 7], tag and NL captions are widely used captioning methods. To allow TIPO model to seamlessly transform captions between different formats or refine captions within the same format, we define specific tasks that enable flexible and precise prompt construction.
5.1 Modeling
TIPO adopts a causal autoregressive LM framework as prompt extension function, $\mathcal{E}$, training to extend and refine user-provided prompts into detailed, content-rich prompts $p_{d}$ from simpler inputs $p_{s}$, thereby effectively balancing specificity with alignment to user intent. This process captures a subset of outputs closely aligned with the user’s original intent, ensuring relevance. Through use of existing datasets for prompt optimization, along with its iterative extension function and application of LM, TIPO achieves significant coverage and diversity in generating high-quality images aligned with training distributions, all without the need for manual prompt engineering.
5.2 Task Definitions
TIPO encompasses three primary tasks: extending tag sequences, extending NL prompts, and generating refined NL prompts. To facilitate these tasks, we define several specific task types and corresponding formatting methods during training:
  1. 1.
    tag_to_long: Use tags as metadata to generate a new NL prompt or extend a user-provided NL prompt.
  2. 2.
    long_to_tag: Use an NL prompt as metadata to extend a tag sequence.
  3. 3.
    short_to_tag: Use the simple prompt $p_{s}$ as metadata to extend a tag sequence.
  4. 4.
    short_to_long: Use a user-input NL prompt as metadata to generate a refined detailed prompt $p_{d}$.
  5. 5.
    short_to_tag_to_long: Use a user-input NL prompt or tag sequence as metadata to generate a refined detailed prompt $p_{d}$.
  6. 6.
    short_to_long_to_tag: Use a user-input NL prompt or generated NL prompt as metadata to extend a tag sequence.
  7. 7.
    tag_to_short_to_long: Use user-input tags or NL prompts as metadata to generate a refined detailed prompt $p_{d}$.
By defining these task types, we can achieve special cases where, for instance, inputting a short description or a short tag sequence results in a full-size tag sequence and NL prompt in a single pass (i.e., generating $T_{d}$ and $p_{d}$ simultaneously) as shown in Section 4.5.
5.3 Model and Training Details
We utilize the LLaMA architecture [46, 47] with models of 200 million and 500 million parameters:
@@ 5.3 Model and Training Details
  • TIPO-200M: Pretrained on the Danbooru2023 [34, 49] and GBC10M [25] datasets for 5 epochs, then fine-tuned on the Danbooru2023, GBC10M, and CoyoHD11M [9] datasets for 3 epochs, resulting in a total of approximately 40 billion non-padding tokens seen.
@@
  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, resulting in a total of approximately 30 billion non-padding tokens seen.
Note that we calculate the token count based on non-padding tokens only. Given the relatively short and variable lengths of data entries, the effective token count may be lower than anticipated.
5.4 Training Procedure
In our experiments, each training pass generates a new prompt of a single type. To cover various prompt types—such as a refined prompt $p_{d}$, an extended NL prompt $S_{d}$, and an extended tag sequence $T_{d}$—the model requires at least three passes. For each dataset entry, we randomly select one of the seven task types to apply and randomly decide how to split the simple prompt from the detailed prompt ($p_{s}$ from $p_{d}$) and the simple tag set from the detailed tag set ($T_{s}$ from $T_{d}$).
This randomization in task selection and prompt splitting effectively increases the training dataset’s size, allowing the model to encounter a broader range of inputs and outputs. This approach enhances the model’s ability to generalize across the various tasks defined in TIPO.
@@ 5.4 Training Procedure
Figure 4: Scenery test examples from 4 different methods.
@@ 6 Evaluation
In this section, we detail the experimental setup used to evaluate the TIPO framework. We outline the formatting conventions for prompts and metadata, define the various tasks employed during training, and describe the models and datasets utilized in our experiments.
@@ 6.1 Experimental Setup
To assess the quality of generated prompts and resulting images, we employed the following evaluation metrics.
FDD (Frechet Dino Distance): Traditionally, the Frechet Inception Distance (FID) [24] has been used to measure the distribution distance between datasets and generated images. However, recent works have shown that FID is not always aligned with human preferences [44]. Therefore, we measure the Frechet Distance using DinoV2 [35] outputs. In our experiments, we used the pretrained ViT-g [18] model from DinoV2 and used the normalized embedding vector to measure the Frechet Distance result.
Aesthetic Score: We compute the Aesthetic Score using the Aesthetic Predictor V2.5 [17]. Higher aesthetic scores indicate images that are perceived as more visually appealing, well-composed, and of higher quality. This metric aims to quantify the subjective visual appeal of generated images, with higher scores suggesting outputs that are more likely to be judged favorably by human viewers.
AI Corrupt Score: The AI Corrupt Score is obtained from the AICorruptMetrics [32] in sdeval [33]. This metric is trained on AI-generated images with human-annotated “corrupt or not” labels. A higher value indicates a higher likelihood of the image being “correct” or “complete”.
All results presented in this section are tested on the TIPO-200M model.
6.2 Scenery Tag Test
In this test, we randomly select 32,768 entries from Danbooru2023 [50, 34] that include the “scenery” tag. We set up the following inputs:
  • “Scenery” + Meta:
    • Retain all metadata categories as described earlier.
    • Only include “scenery” as the content tag.
  • “Scenery” + Meta + TIPO:
    • Use “scenery” + meta as input.
    • Extend $T_{s}$ (scenery) to $T_{d}$ and generate $p_{d}$ from $T_{d}$.
For the scenery tag test, we generate one image from each prompt using the Kohaku-XL-zeta [^2] https://huggingface.co/KBlueLeaf/Kohaku-XL-Zeta model, which is a fine-tuned SDXL [37] model on the Danbooru dataset [34].
Figure 4 illustrates the comparative visual result by various methods. The results reveal significant biases in Prompt DB-based [21] and RL-based [22] approaches. The former consistently generates snowy mountain scenes, while the latter produces near-future fantasy scenery, often resulting in corrupted images. The Direct LLM method, utilizing GPT-4o [2], shows improved results but still exhibits notable bias; out of 9 generated prompts, 8 exclusively describe natural scenes only without any human-made structures.
In contrast, TIPO demonstrates superior performance in reproducing the dataset’s distribution. While primarily generating nature scenes, it also incorporates architectural elements and artifacts, more accurately reflecting the diverse content of the original dataset. This balanced output suggests that TIPO effectively mitigates the biases observed in other methods.

### v2

@@ Abstract
TIPO (Text to Image with text pre-sampling for Prompt Optimization) is an innovative framework designed to enhance text-to-image (T2I) generation by language model (LM) for automatic prompt engineering. By refining and extending user-provided prompts, TIPO bridges the gap between simple inputs and the detailed prompts required for high-quality image generation. Unlike previous approaches that rely on Large Language Models (LLMs) or reinforcement learning (RL), TIPO adjusts user input prompts with the distribution of a trained prompt dataset, eliminating the need for complex runtime cost via lightweight model. This pre-sampling approach enables efficient and scalable prompt optimization. Experimental results demonstrate TIPO’s effectiveness in improving aesthetic scores, reducing image corruption, and better aligning generated images with dataset distributions. These findings highlight the critical role of prompt engineering in T2I systems and open avenues for broader applications of automatic prompt refinement.
@@ 1 Introduction
Recent advancements in Text-to-Image (T2I) generative models have revolutionized creative applications [47, 7, 21, 55, 51, 52, 57, 53, 48, 56, 15, 14, 34, 22, 8]. These models tend to perform better when provided with longer, more detailed prompts, which describes specific elements like style, composition, and context. However, this reliance on highly descriptive inputs is a significant bottleneck for generating high-quality images, especially for users who prefer or require simpler inputs. The need for intricate, nuanced prompts often results in a higher barrier to entry, limiting accessibility and ease of use for those unfamiliar with the complexities of prompt engineering.
Previous works have attempted to extend prompts with optimization through various methods. Some approaches use LLMs to expand prompts [41]; however, due to the misalignment between the LLM’s text-rich training corpus and the T2I model’s caption-focused corpus, only a limited set of manually-engineered prompts yield effective results. Other methods rely on curated prompt sets [5, 17], which, while potentially incorporating user preferences, fail to fully leverage the T2I model’s trained prompt distribution. Additionally, reinforcement learning-based techniques [26] require multi-turn inference, incurring high computational costs, and often struggle to generate diverse outputs due to the model’s inherent corpus limitations.
@@
Figure 1: Comparison of prompt optimization methods. (a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not T2I model. (b) relies on a curated prompt database, enhancing detail but limiting variety by not fully leveraging the T2I model’s learned distribution. (c) optimizes using the scorer with RL, requiring multi-turn inference with additional cost. (d) aligns prompts with the T2I model’s training distribution, ensuring detailed and various topic-related prompt generation that fits the target T2I model.
This paper presents TIPO’s theoretical foundations (Section 3), implementation details (Section 4), and experimental results (Section 5) demonstrating its effectiveness in improving aesthetic scores [19], reducing image corruption [43, 44], and better aligning generated images with dataset distributions [28, 46]. Our findings highlight the critical role of prompt optimization in advancing T2I technology and open avenues for broader applications of automatic prompt refinement.
@@ 2.3 Image Output-Based Tuning
[…]
@@ 3.1 Motivation and Challenges
Nowadays text-conditioned models [8, 48, 21] generate outputs based on user-provided text inputs. However, when users seek high-quality outputs or wish to explore diverse outputs while preserving key aspects of their input, they often need to manually expand their prompts. This manual prompt expansion can be time-consuming and labor-intensive.
To address these challenges, previous works have aimed to automate the prompt optimization process. Some methods attempt to expand prompts using the familiar vocabulary and stylistic tendencies of the model [25], while others rely on direct feedback from the model to iteratively improve prompt alignment [26]. However, these approaches have limitations: if the structure or language does not fully align with the model’s prompting conventions, performance may degrade, and feedback-based approaches often involve significant computational costs.
Therefore, effective prompt optimization for text-conditioned models requires accurately capturing user intent and expanding the prompt in a way that is both precise and aligned with the model’s conventions for interpreting text input. This need for precision and alignment motivates a systematic approach to prompt formulation, which we define in the following section.
@@ 3.2 Formulation
The simplest approach to extend a simple prompt is by adding random words. However, this approach has limitations, as it often introduces internally conflicting information that results in out-of-distribution (OOD) prompt inputs, leading the model to generate nonsensical images or the model to generate not user intent images. In such cases, although each added word may be meaningful in isolation, these random combinations can create prompts that are internally contradictory, as shown in Figure 2 (c). For example, the generated images may deviate from the user’s intent by not depicting an astronaut (left top), producing a horse-shaped tank (right top) instead of a horse, omitting Mars (left bottom), or even including a random snake (right bottom), which introduces nonsensical elements. To represent this random extension method, let $p_{r_{i}}$ denote the $i^{th}$ prompt created by adding random words to $p_{s}$. The resulting union of distributions fails to approximate the desired distribution.
@@
This scenario is illustrated in Figure 2 (c), where “Simple Prompt + Random Words” shows a pointy distribution extending beyond the bounds of the desired distribution or even the “image” distribution itself.
4 TIPO Framework and Model
@@ 4.1 Distribution-Aligned Prompt Extension
where $p_{t_{i}}$ denotes the $i^{th}$ TIPO-generated prompt derived from $p_{s}$, and $\mathcal{I}_{p_{s}}$ represents the target distribution of image outputs associated with $p_{s}$. Through this systematic prompt refinement, Figure 2 (d) “Simple Prompt + TIPO” illustrates the broader image space exploration enabled by TIPO while retaining relevance to the original prompt, effectively balancing diversity with coherence.
@@ 4.3 Prompt Formatting
4.4 Task Definitions
Figure 3: TIPO’s task flow within a single generation cycle. Starting from any node, each arrow represents a sequential extension step, with other prompt used as metadata if provided. This task design enables efficient and flexible prompt extension across multiple tasks.
TIPO defines three core tasks: extending tag sequences, extending NL prompts, and generating refined NL prompts. Each task is represented by an arrow in Figure 3, illustrating how TIPO can use tag and NL prompts interchangeably as input sources to generate detailed outputs within a single generation cycle. When an arrow cycle is established, TIPO can complete multiple extension within a single generation pass.
For example, TIPO can start from a short NL prompt to produce a detailed tag sequence, or it can combine tasks like short_to_tag and tag_to_long to create a composite task, short_to_tag_to_long, which generates an extended NL prompt from an initial short NL input via tag generation. Further details on specific task sequences are provided in Appendix.
@@ 4.5 An Example Scenario and Workflow
Figure 4: An example scenario of TIPO workflow. (a) A generated image and prompts. (b) Prompt optimization of TIPO, starting from simple user input $p_{s}$ to detailed final output $p_{d}$. The shading from gray to light sky blue represents an increase in context richness in the prompt.
@@
$\begin{array}[]{ll}S_{s}&=\{\textsf{A young girl with long hair...}\},\quad\verb|<meta>|=\emptyset{}\\ T_{s}&=\left\{\textsf{outdoors, scenery, water, wind, landscape, ...}\right\}\end{array}$
@@
    short_to_tag_to_long: With both $T_{d}$ and $S_{s}$ in the metadata, TIPO generates a comprehensive long NL prompt $S_{d}$, ensuring a more detailed prompt.
@@
This iterative process allows TIPO to build prompts that are both detailed and contextually aligned with user inputs, through short_to_tag, tag_to_long, and short_to_tag_to_long tasks within a single generation cycle.
4.6 Model Architecture
TIPO utilizes a causal autoregressive language model framework as its prompt extension function, $\mathcal{E}$, to iteratively extend and refine user-provided prompts into more detailed, content-rich prompts $p_{d}$ based on simpler inputs $p_{s}$. Our implementation leverages the LLaMA architecture [61, 62, 3], employing models with 200M and 500M parameters.
@@ 4.6 Model Architecture
  • TIPO-200M: Pretrained on the Danbooru2023 [45, 65] and GBC10M [29] datasets for 5 epochs, then fine-tuned with Danbooru2023, GBC10M, and CoyoHD11M [12] for 3 epochs, covering approximately 40 billion tokens.
@@
  • TIPO-500M: Pretrained on the Danbooru2023, GBC10M, and CoyoHD11M datasets for 5 epochs, covering approximately 30 billion tokens.
Throughout training, non-padding tokens were used for token count calculations, with the effective count reflecting the relatively short and variable length of data entries.
4.7 Training Process
In each training pass, TIPO applies dataset augmentation by selecting tasks from all possible task combinations such as short_to_tag_to_long, including extending tag sequences, NL prompt extension, and refined NL prompt generation. This random selection across all task combinations significantly broadens the dataset’s effective scope, enhancing model generalization without additional manual engineering or runtime tuning.
@@ 4.7 Training Process
Figure 5: In-domain test regard to scenery test examples from 4 different methods.
(a) Scenery Tag Only
(b) Scenery Tag + TIPO
(c) Truncated Long Prompt
(d) Truncated Prompt + TIPO
Figure 6: Comparison of generated images: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated long prompt, and (d) TIPO-enhanced truncated prompt. TIPO adds detail and maintains variety, yielding coherent images even from minimal prompts.
@@ 5.1 Experimental Setup
In this section, we detail the experimental setup used to evaluate prompt optimization framework. To assess the quality of generated prompts and resulting images, we employed the following evaluation metrics.
FDD (Frechet Dino Distance).
Traditionally, the Frechet Inception Distance (FID) [28] has been used to measure the distribution distance between datasets and generated images. However, recent works have shown that FID is not always aligned with human preferences [58]. Therefore, we measure the Frechet Distance using DinoV2 [46] outputs. In our experiments, we used the pretrained ViT-g [20] model from DinoV2 and used the normalized embedding vector to measure the Frechet Distance result.
Aesthetic Score.
We compute the Aesthetic Score using the Aesthetic Predictor V2.5 [19]. Higher aesthetic scores indicate images that are perceived as more visually appealing, well-composed, and of higher quality. This metric aims to quantify the subjective visual appeal of generated images, with higher scores suggesting outputs that are more likely to be judged favorably by human viewers.
AI Corrupt Score.
The AI Corrupt Score is obtained from the AICorruptMetrics [43] in sdeval [44]. This metric is trained on AI-generated images with human-annotated “corrupt or not” labels. A higher value indicates a higher likelihood of the image being “correct” or “complete”.
Vendi Score.
To quantify the diversity of generated images, we employ the Vendi Score metric based on DinoV2 [46] embeddings. The computation involves calculating a similarity matrix $K$ using cosine similarity between image embeddings, followed by computing the von Neumann entropy of the normalized matrix $K/n$, adhering to the standard Vendi Score methodology. This metric provides a quantitative measure of the distributional diversity within a set of generated images. Higher Vendi Scores indicate greater inter-image diversity, reflecting the model’s capability to generate varied and distinct outputs.
Extension Functions.
We used both GPT-4o-mini [2] and LLaMA-3.2-1B [3] as Direct LLM models for prompt generation, balancing performance and efficiency. GPT-4o-mini, as a lightweight closed-source model, offers speed and affordability, while LLaMA-3.2-1B represents a similar scale to TIPO in open-source models. Also, Prompt DB [17] and Promptist [26] serve as baseline methods for prompt extension.
5.2 In-domain test regard to scenery tag
To evaluate the model’s capability to generate diverse and accurate outputs, we chose the scenery tag, which includes both natural landscapes and human-made structures. In this test, we randomly select 32,768 entries from Danbooru2023 [65, 45] that include the “scenery” tag. We set up the following prompt:
<meta> + scenery + $\mathcal{E}$>
Then, we apply extension functions to extend only the scenery, $T_{s}$, as the <Content> tag. We generate one image from each prompt using the Kohaku-XL-zeta [^2] https://huggingface.co/KBlueLeaf/Kohaku-XL-Zeta model, which is a fine-tuned SDXL [48] model on the Danbooru dataset [45].